# Segment Any 3D Gaussians

### v2

@@ 3.4 Inference
With well-trained Gaussian affinity features, SAGA can conduct various segmentation tasks in the 3D space. For promptable segmentation, SAGA takes 2D point prompts at specific view and the scale as input. Then, SAGA segments the 3D target by matching scale-gated 3D Gaussian affinity features with the 2D query features selected from the rendered feature map according to prompt points. For automatic scene decomposing, SAGA employs HDBSCAN to cluster the affinity features directly in the 3D space [^2] For efficiency, SAGA uniformly selects 1% of the Gaussians from the 3D-GS model for clustering.. Additionally, we design a vote-based segmentation mechanism to integrate SAGA with CLIP [42] for conducting open vocabulary segmentation (see Section A.3).
@@ 4 Experiments
In this section, we demonstrate the effectiveness of SAGA both quantitatively and qualitatively. For implementation details and specific experiment settings, please refer to Section A.2.
@@ 4.1 Datasets
For promptable segmentation experiments, we utilize two datasets: Neural Volumetric Object Selection (NVOS)[43] and SPIn-NeRF[38]. The former is derived from the LLFF dataset [36] and the latter is a combination of subsets of data from established NeRF-related datasets [36, 37, 28, 25, 13]. For open-vocabulary segmentation experiments, we adopt the 3D-OVS dataset [30]. For qualitative analysis, we employ various datasets including LLFF [36], MIP-360 [1], Tanks&Temple [25], and Replica [47]. These datasets encompass indoor and outdoor scenes, forward-facing and 360-degree scenes, as well as synthetic and real scenes.
@@ 4.2.1 Promptable Segmentation
SPIn-NeRF Dataset
Results on the SPIn-NeRF dataset can be found in Table 2. SA3D performs on par with the previous SOTA OmniSeg3D. The minor performance degradation is attributed to sub-optimal geometry learned by 3D-GS. Specifically, in the “fork” scene, which features a metal fork on a table, the segmentation target is simple (i.e., the fork). However, 3D-GS models the reflection effects of the fork with numerous outlier Gaussians that are not aligned with the exact geometry. Excluding these Gaussians results in empty holes in the segmentation mask for certain views, while including them introduces noise in other views. Nevertheless, we believe the segmentation accuracy of SAGA can meet most requirements.
[…]
@@
NVOS [43] 70.1 92.0
ISRF [15] 83.8 96.4
SA3D [7] 90.3 98.2
OmniSeg3D [53] 91.7 98.4
SA3D-GS [6] 92.2 98.5
@@
[…]
@@
MVSeg [38] 90.9 98.9
SA3D [7] 92.4 98.9
OmniSeg3D [53] 94.3 99.3
SA-GS [20] 89.9 98.7
SA3D-GS [6] 93.2 99.1
@@
Figure 4: Ablation study on effects of local feature smoothing (Smooth) and feature norm regularization (Feature Norm). Outliers are primarily eliminated through local feature smoothing. Feature norm regularization helps features of inner Gaussians align better with those of the surface Gaussians.
4.2.2 Open-Vocabulary Semantic Segmentation
To conduct open-vocabulary semantic segmentation, we follow the setting of 3D-OVS where all categories are input to CLIP to obtain their textual features. These features are then used to vote for all clusters as introduced in Section 3.4. For each cluster, the category with the highest score is assigned as its label. As shown in Table 3, SAGA demonstrates superior results across all scenes in the 3D-OVS dataset. However, we acknowledge that this result does not imply that SAGA consistently outperforms previous open-vocabulary radiance field segmentation methods. Please refer to Section A.3 for a detailed discussion on the limitations of SAGA in open-vocabulary segmentation.
Table 3: Open vocabulary segmentation results on 3D-OVS dataset (mIoU).
@@ 4.2.2 Open-Vocabulary Semantic Segmentation
LERF [22] 73.5 53.2 46.6 27.0 73.7 54.8
3D-OVS [30] 89.5 89.3 92.8 74.0 88.2 86.8
LangSplat [41] 92.5 94.2 94.1 90.0 96.1 93.4
N2F2 [41] 93.8 92.6 93.5 92.1 96.3 93.9
SAGA (ours) 97.4 95.4 96.8 93.5 96.6 96.0
4.2.3 Time Consumption Analysis
@@
SA3D [7] - 45 s
SA-GS [20] - 15 s
ISRF [15] 2.5 mins 3.3 s
OmniSeg3D [53] 15~40 mins 50~100 ms
GARField [23] 20~60 mins 30~70 ms
@@
In Table 4, we reveal the time consumption of SAGA and compare it with existing promptable segmentation methods in radiance fields. ISRF trains a feature field by mimicking the 2D multi-view visual features extracted by DINO [5], thus enjoys faster convergence speed. However, this approach results in less accurate segmentation, necessitating extensive post-processing. SA3D and SA-GS employ an iterative mask refinement pipeline, which eliminates the need for training but incurs significant inference time consumption. Compared to methods that distill segmentation capabilities from SAM masks, such as OmniSeg3D and GARField, SAGA demonstrates much faster inference speed and comparable training speed.
@@ 4.3 Qualitative Results
As shown in Figure 3, SAGA can achieve fine-grained segmentation at different scales in various types of scenes. The learned Gaussian affinity features are compact enough for scene decomposition through simple clustering across the entire 3D-GS model. Note that there are some needle-like burrs on the edges of the segmented targets. This artifact is due to the nature of 3D-GS, which aims to fit multi-view RGB color given camera poses without object awareness. For more visualization results, please refer to our appendix.
@@ 4.4 Ablation Study
Local Feature Smoothing & Feature Norm Regularization
Both local feature smoothing and feature norm regularization impose constraints on the Gaussian affinity features. We present visualization results to illustrate their roles intuitively.
As shown in Figure 4, when segmenting a 3D object with a cosine similarity threshold of 0.75, SAGA without the feature smoothing mechanism shows many false positives. This indicates This indicates that outliers are primarily eliminated by the feature smoothing operation.
Unlike local feature smoothing, feature norm regularization primarily impacts the Gaussians within objects. When raising the similarity score threshold to 0.95, the apple segmented by SAGA with feature norm regularization remains intact, while the one without it quickly becomes translucent. This phenomenon supports our assumption that 3D features are not perfectly aligned with 2D features, as introduced in Section 3.3.2. Imposing feature norm regularization helps align the affinity features of 3D Gaussians by pulling the features along a ray in the same direction.
@@ 5 Limitation
SAGA learns the affinity features from multi-view 2D masks extracted by SAM. This makes SAGA hardly segment objects that are not appeared in these masks. Enhancing the generalization ability of SAGA to unrecognized targets during the automatic extraction stage is a promising direction.
@@ 6 Conclusion
In this paper, we propose SAGA, a 3D promptable segmentation method for 3D Gaussian Splatting (3D-GS). SAGA injects the segmentation capability of SAM into Gaussian affinity features for all 3D Gaussians in the 3D-GS model, endowing them with a new property towards segmentation. To simultaneously preserve the multi-granularity segmentation ability of SAM and the efficiency of 3D-GS, SAGA introduces a lightweight scale-gate mechanism, which adapts the affinity features according to different 3D physical scales with minimal computation overhead. After training, SAGA can achieve fine-grained 3D segmentation within milliseconds. Comprehensive experiments are conducted to demonstrate the effectiveness of SAGA. As one of the first methods addressing promptable segmentation in 3D-GS, the simplicity and effectiveness of SAGA pave the way for future advancements in this field.
@@ Appendix A Appendix
In this appendix we provide the concrete training strategy of SAGA and implementation details. Then, we provide details about the open-vocabulary segmentation ability of SAGA and analyze its limitation. We also provide more visualization results to show the effectiveness of SAGA intuitively.
@@ A.1 Detailed Additional Training Stategy
$\mathcal{L}=$ $\frac{1}{|\mathcal{Q}^{\texttt{in}}\cup\phi(\mathcal{Q}^{\texttt{pos}})|}\sum_{(\mathbf{p}_{1},\mathbf{p}_{2})\in\mathcal{Q}^{\texttt{in}}\cup\phi(\mathcal{Q}^{\texttt{pos}})}\omega(\mathbf{p}_{1},\mathbf{p}_{2})\mathcal{L}_{\texttt{corr}}(\mathbf{p}_{1},\mathbf{p}_{2})+$ (13)
$\frac{1}{|\mathcal{Q}^{\texttt{in}}\cup\phi(\mathcal{Q}^{\texttt{neg}})|}\sum_{(\mathbf{p}_{1},\mathbf{p}_{2})\in\mathcal{Q}^{\texttt{in}}\cup\phi(\mathcal{Q}^{\texttt{neg}})}\omega(\mathbf{p}_{1},\mathbf{p}_{2})\mathcal{L}_{\texttt{corr}}(\mathbf{p}_{1},\mathbf{p}_{2})+$
@@ A.3 Vote-based Open-vocabulary Segmentation
To enable open-vocabulary 3D segmentation in radiance fields, previous studies [22, 30, 41, 2, 27] focus on aligning 3D language feature fields with the visual features extracted by CLIP [42] image encoder. Then 3D segmentation can be achieved by querying the language fields with the textual features. The process of training language feature field can be regarded as multi-view feature fusing, which, in essence, is a kind of vote mechanism. This motivate us to see whether SAGA can handle open-vocabulary segmentation with well-trained Gaussian affinity features at minimum modification.
Readers may wonder why we do not use grounding methods like Grounding-DINO [32] to locate the object and convert the location into point prompts, which can then be fed to SAGA. To answer this question, we emphasize the setting of open-vocabulary segmentation. We follow a stricter approach than previous studies, such as SA3D [7] and GaussianGrouping [52], which rely on Grounded-SAM to convert language prompts to visual prompts in a specific view. Specifying this view introduces additional prior knowledge. In SAGA, users do not need to query a specific view; all that is required for open-vocabulary segmentation is a text prompt. With this idea in mind, we introduce a vote-based open-vocabulary segmentation strategy.
@@
For each 2D mask $\mathbf{M}_{\mathbf{I}}$ of image $\mathbf{I}$, we extract its visual feature by feeding the masked image $\mathbf{I}\odot\mathbf{M}_{\mathbf{I}}$ to the CLIP visual encoder. During inference, given any text prompt, a relevancy score $r_{\mathbf{M}}$ is assigned to each mask $\mathbf{M}$ by comparing the textual feature with the visual feature of the mask [^4] We adopt the relevancy score introduced in LERF [22], which has been proven robust.. For a cluster centroid $\mathbf{T}$ in $\mathcal{V}$, the relevancy scores of its corresponding 2D masks are aggregated to form its final relevancy score, i.e.,
@@
The vote-based open-vocabulary segmentation strategy encounters difficulties in certain scenarios. For instance, when considering a bowl of noodles with an egg in it, SAGA ideally should distinguish between categories such as the egg, the egg with noodles (contents of the bowl), and the bowl with noodles and egg. However, because "egg" is included in masks at larger scales, CLIP often misclassifies larger objects as "egg." This issue is inherently rooted in the multi-granularity ambiguity in semantics. Addressing this problem is a promising research direction.
@@
Another limitation is common among current CLIP-SAM based methods, such as LangSplat [41] and N2F2 [2]. Both SAGA and these methods first use SAM to segment the images and then use CLIP to attach semantics to these segments by segmenting the corresponding regions of the image and feeding them to CLIP. However, the effectiveness of the CLIP visual encoder also depends on context. For example, SAM sometimes segments the texture on a wall. Without seeing the entire wall, CLIP struggles to recognize it. This lack of context hinders the ability to ground the segments accurately. This is an important yet currently unexplored issue.
A.4 SAGA with Hybrid / Implicit Radiance Fields

### v3

@@ 3.5 Inference
With well-trained Gaussian affinity features, SAGA can conduct various segmentation tasks in the 3D space. For promptable segmentation, SAGA takes 2D point prompts at specific view and the scale as input. Then, SAGA segments the 3D target by matching scale-gated 3D Gaussian affinity features with the 2D query features selected from the rendered feature map according to prompt points. For automatic scene decomposing, SAGA employs HDBSCAN to cluster the affinity features directly in the 3D space [^2] For efficiency, SAGA uniformly selects 1% of the Gaussians from the 3D-GS model for clustering.. Additionally, we design a vote-based segmentation mechanism to integrate SAGA with CLIP for conducting open vocabulary segmentation (see Section A.3 of the supplement).
@@ 4 Experiments
In this section, we demonstrate the effectiveness of SAGA both quantitatively and qualitatively. For implementation details and experiment settings, please refer to Section A.2.
@@ 4.1 Datasets
For promptable segmentation experiments, we utilize two datasets: NVOS (Ren et al. 2022) and SPIn-NeRF(Mirzaei et al. 2023). The former is derived from the LLFF dataset (Mildenhall et al. 2019) and the latter is a combination of subsets of data from established NeRF-related datasets (Mildenhall et al. 2019, 2020; Lin et al. 2022; Knapitsch et al. 2017; Fridovich-Keil et al. 2022). For open-vocabulary segmentation experiments, we adopt the 3D-OVS dataset (Liu et al. 2023a). For qualitative analysis (Figure 3 and 4), we employ various datasets including LLFF (Mildenhall et al. 2019), MIP-360 (Barron et al. 2022), Tanks&Temple (Knapitsch et al. 2017), and Replica (Straub et al. 2019). These datasets encompass indoor and outdoor scenes, forward-facing and 360-degree scenes, as well as synthetic and real scenes.
@@ 4.2 Quantitative Results
NVOS (Ren et al. 2022) 70.1 92.0
ISRF (Goel et al. 2023) 83.8 96.4
SA3D (Cen et al. 2023) 90.3 98.2
OmniSeg3D (Ying et al. 2024) 91.7 98.4
GauGroup (Ye et al. 2024) 85.6 97.3
SA3D-GS (Cen et al. 2024) 92.2 98.5
@@
SPIn-NeRF Dataset
Results on the SPIn-NeRF dataset can be found in Table 2. SA3D performs on par with the previous SOTA OmniSeg3D. The minor performance degradation is attributed to sub-optimal geometry learned by 3D-GS. For example, 3D-GS models the reflection effects with numerous outlier Gaussians that are not aligned with the exact geometry of the object. Excluding these Gaussians results in empty holes in the segmentation mask for certain views, while including them introduces noise in other views. Nevertheless, we believe the segmentation accuracy of SAGA can meet most requirements.
@@
MVSeg (Mirzaei et al. 2023) 90.9 98.9
SA3D (Cen et al. 2023) 92.4 98.9
OmniSeg3D (Ying et al. 2024) 94.3 99.3
GauGroup (Ye et al. 2024) 86.5 98.9
SA-GS (Hu et al. 2024) 89.9 98.7
SA3D-GS (Cen et al. 2024) 93.2 99.1
@@
Open-Vocabulary Semantic Segmentation
As shown in Table 3, SAGA demonstrates superior results across all scenes in the 3D-OVS dataset. For more details about open-vocabulary segmentation, please refer to Section A.3.
@@
LERF 73.5 53.2 46.6 27.0 73.7 54.8
3D-OVS 89.5 89.3 92.8 74.0 88.2 86.8
LangSplat 92.5 94.2 94.1 90.0 96.1 93.4
N2F2 93.8 92.6 93.5 92.1 96.3 93.9
SAGA 97.4 95.4 96.8 93.5 96.6 96.0
Table 3: Results on 3D-OVS dataset (mIoU).
Time Consumption Analysis
In Table 4, we reveal the time consumption of SAGA and compare it with existing promptable segmentation methods in radiance fields. ISRF trains a feature field by mimicking the 2D multi-view visual features extracted by DINO (Caron et al. 2021), thus enjoys faster convergence speed. However, this leads to less accurate segmentation, necessitating extensive post-processing. SA3D and SA-GS employ an iterative mask refinement pipeline, which eliminates the need for training but incurs significant inference time consumption. Compared to methods that distill segmentation capabilities from SAM masks, such as OmniSeg3D and GARField, SAGA demonstrates much faster inference speed and comparable training speed.
@@
SA3D (Cen et al. 2023) - 45 s
SA-GS (Hu et al. 2024) - 15 s
ISRF (Goel et al. 2023) 2.5 mins 3.3 s
OmniSeg3D (Ying et al. 2024) 15~40 mins 50~100 ms
GARField (Kim et al. 2024) 20~60 mins 30~70 ms
@@ 4.3 Qualitative Results
Figure 3 shows that SAGA achieves fine-grained segmentation at various scales across different scenes. The compact Gaussian affinity features enable scene decomposition through simple clustering in the 3D-GS model. The needle-like artifacts on the edges of segmented targets are due to 3D-GS overfitting multi-view RGB without object awareness. Building upon 3D-GS, which can capture high-frequency texture details, SAGA can effectively segment thin, fine-grained structures, as shown in Figure 4. By shrinking the 3D Gaussians, we reveal the underlying structural modeling capabilities of 3D-GS and demonstrate the completeness of the segmentation results. Since GARField lacks quantitative results on the NVOS and SPIn-NeRF datasets, we provide qualitative comparisons in Section A.6 to highlight the effectiveness of SAGA’s learned features.
Figure 4: SAGA can maintain the high frequency texture details captured by 3D-GS. We reveal the inherent structure of these details by shrinking the Gaussians by 60%.
@@ 4.4 Ablation Study
Local Feature Smoothing (LFS) & Feature Norm Regularization (FNR)
Both LFS and FNR impose constraints on the Gaussian affinity features. We present visualization results to illustrate their roles.
In Figure 5, when segmenting a 3D object with a cosine similarity threshold of 0.75, without the feature smoothing, the result shows many false positives, which reveals that outliers are primarily eliminated by the feature smoothing operation. Unlike local feature smoothing, feature norm regularization primarily impacts the Gaussians within objects. When raising the similarity score threshold to 0.95, the apple segmented by SAGA with feature norm regularization remains intact, while the one without it quickly becomes translucent. This phenomenon supports our assumption that 3D features are not perfectly aligned with 2D features, as introduced in Section 3.4. Imposing feature norm regularization helps align the affinity features of 3D Gaussians by pulling the features along a ray in the same direction.
Figure 5: Ablation study on effects of local feature smoothing (Smooth) and feature norm regularization (Feature Norm). Outliers are primarily eliminated through local feature smoothing. Feature norm regularization helps features of inner Gaussians align better with those of the surface.
Figure 6: Failure cases of SAGA. The targets of interest are labeled by red border.
@@ 5 Limitation
SAGA learns the affinity features from multi-view 2D masks extracted by SAM. This makes SAGA hardly segment objects that are not appeared in these masks. As shown in Figure 6, this limitation is particularly evident when the target of interest is small. Enhancing the generalization ability of SAGA to unrecognized targets during the automatic extraction stage is a promising direction.
@@ 6 Conclusion
In this paper, we propose SAGA, a 3D promptable segmentation method for 3D Gaussian Splatting (3D-GS). SAGA injects the segmentation capability of SAM into Gaussian affinity features for all 3D Gaussians in a 3D-GS model, endowing them with a new property towards segmentation. To preserve the multi-granularity segmentation ability of SAM and the efficiency of 3D-GS, SAGA introduces a lightweight scale-gate mechanism, which adapts the affinity features according to different 3D physical scales with minimal computation overhead. After training, SAGA can achieve real-time fine-grained 3D segmentation. Comprehensive experiments are conducted to demonstrate the effectiveness of SAGA. As one of the first methods addressing promptable segmentation in 3D-GS, the simplicity and effectiveness of SAGA pave the way for future advancements in this field.
Acknowledgments
This work was supported by NSFC 62322604, NSFC 62176159, Shanghai Municipal Science and Technology Major Project 2021SHZDZX0102.
@@ Appendix A Appendix
Figure A1: Adapting the scale gate mechanism with GARField achieves competitive results, demonstrating the potential of SAGA across different radiance fields.
In this appendix we provide the concrete training strategy (Section A.1) of SAGA and implementation details (Section A.2). Then, we provide details about the open-vocabulary segmentation ability of SAGA and analyze its limitation (Section A.3). We also provide an interpretability analysis about the scale gate mechanism (Section A.4) to reveal the underlying principle of SAGA. Section A.5 evaluates the robustness and generalizability of SAGA by applying it to more kinds of radiance fields, and Section A.6 presents additional visualization results to demonstrate its effectiveness.
@@ A.1 Detailed Additional Training Stategy
$\mathcal{L}=$ $\frac{\sum_{(\mathbf{p}_{1},\mathbf{p}_{2})\in\mathcal{Q}^{\texttt{in}}\cup\phi(\mathcal{Q}^{\texttt{pos}})}\omega(\mathbf{p}_{1},\mathbf{p}_{2})\mathcal{L}_{\texttt{corr}}(\mathbf{p}_{1},\mathbf{p}_{2})}{|\mathcal{Q}^{\texttt{in}}\cup\phi(\mathcal{Q}^{\texttt{pos}})|}+$ (13)
$\frac{\sum_{(\mathbf{p}_{1},\mathbf{p}_{2})\in\mathcal{Q}^{\texttt{in}}\cup\phi(\mathcal{Q}^{\texttt{neg}})}\omega(\mathbf{p}_{1},\mathbf{p}_{2})\mathcal{L}_{\texttt{corr}}(\mathbf{p}_{1},\mathbf{p}_{2})}{|\mathcal{Q}^{\texttt{in}}\cup\phi(\mathcal{Q}^{\texttt{neg}})|}+$
@@ A.3 Vote-based Open-vocabulary Segmentation
To enable open-vocabulary 3D segmentation in radiance fields, previous studies (Kerr et al. 2023; Liu et al. 2023a; Qin et al. 2024; Bhalgat et al. 2024; Liao et al. 2024) focus on aligning 3D language feature fields with the visual features extracted by CLIP (Radford et al. 2021) image encoder. Then 3D segmentation can be achieved by querying the language fields with the textual features. The process of training language feature field can be regarded as multi-view feature fusing, which, in essence, is a kind of vote mechanism. This motivate us to see whether SAGA can handle open-vocabulary segmentation with well-trained Gaussian affinity features at minimum modification.
Readers may wonder why we do not use grounding methods like Grounding-DINO (Liu et al. 2024a) to locate the object and convert the location into point prompts, which can then be fed to SAGA. To answer this question, we emphasize the setting of open-vocabulary segmentation. We follow a stricter approach than previous studies, such as SA3D (Cen et al. 2023) and GaussianGrouping (Ye et al. 2024), which rely on Grounded-SAM to convert language prompts to visual prompts in a specific view. Specifying this view introduces additional prior knowledge. In SAGA, users do not need to query a specific view; all that is required for open-vocabulary segmentation is a text prompt. With this idea in mind, we introduce a vote-based open-vocabulary segmentation strategy.
@@
For each 2D mask $\mathbf{M}_{\mathbf{I}}$ of image $\mathbf{I}$, we extract its visual feature by feeding the masked image $\mathbf{I}\odot\mathbf{M}_{\mathbf{I}}$ to the CLIP visual encoder. During inference, given any text prompt, a relevancy score $r_{\mathbf{M}}$ is assigned to each mask $\mathbf{M}$ by comparing the textual feature with the visual feature of the mask [^4] We adopt the relevancy score introduced in LERF (Kerr et al. 2023), which has been proven robust.. For a cluster centroid $\mathbf{T}$ in $\mathcal{V}$, the relevancy scores of its corresponding 2D masks are aggregated to form its final relevancy score, i.e.,
@@
The vote-based open-vocabulary segmentation strategy encounters difficulties in certain scenarios. For instance, when considering a bowl of noodles with an egg in it, SAGA ideally should distinguish between categories such as the egg, the egg with noodles (contents of the bowl), and the bowl with noodles and egg. However, because ”egg” is included in masks at larger scales, CLIP often misclassifies larger objects as ”egg.” This issue is inherently rooted in the multi-granularity ambiguity in semantics. Addressing this problem is a promising research direction.
@@
Another limitation is common among current CLIP-SAM based methods, such as LangSplat (Qin et al. 2024) and N2F2 (Bhalgat et al. 2024). Both SAGA and these methods first use SAM to segment the images and then use CLIP to attach semantics to these segments by segmenting the corresponding regions of the image and feeding them to CLIP. However, the effectiveness of the CLIP visual encoder also depends on context. For example, SAM sometimes segments the texture on a wall. Without seeing the entire wall, CLIP struggles to recognize it. This lack of context hinders the ability to ground the segments accurately. This is an important yet currently unexplored issue.
A.4 Interpretability Analysis
To better understand the scale gate mechanism, we analyze the weights of the learned scale gates through a statistical analysis across 47 scenes, each containing $32$ scale gates, resulting in a total of $47\times 32$ entries. Within these entries, 36.1% (543) of the gates are positive, while 63.9% (961) are negative, meaning that a typical scene has about 12 positive gates and 20 negative gates. In other words, when larger-scale features are used as input, more gates tend to close, and fewer gates remain open, which aligns with the intuitive understanding that finer-grained segmentation requires more features to effectively capture detailed information.
A.5 SAGA with Hybrid Radiance Fields